OpenStudy (anonymous):

HELP?! At his son's birth, a man invested $2,000 in savings at 6% for his son's college education. Approximately how much, to the nearest dollar, will be available in 19 years? Answer is $6051.20 Because of inflation, money is depreciating in real value. If the rate of inflation is 8% a year, what is the current real value, to the nearest dollar, from the preceding problem above? After 19 years, did the father and son have more or less money, as measured by current real value, than they did at the son's birth? WHAT IS THE REAL VALUE?

OpenStudy (anonymous):

is the investment compounded monthly?

OpenStudy (anonymous):

I'm not sure, this is all they provided me with @DemolisionWolf

OpenStudy (anonymous):

sorry, i failed to read the whole question. I get what is being asked now. "what is the current real value, to the nearest dollar, from the preceding problem above?" 'prededing problem above' value is 6051.20, which grew over 19 years at 6% starting from 2000 dollars. if interest is 8% each year, and the interest we were using was only 6%, then we are actually loosing money each year when investing at 6%. But lets pretend we invested at 8% (the same value that inflation is), then we wouldn't actually make any money, nor would we loose any money. making sense so far?

OpenStudy (anonymous):

"what is the current real value, to the nearest dollar, from the preceding problem above?" so, we have a 'future' value of 6051.20, we need to figure out what that is worth today considering inflation of 8%. would it be 2000 dollars still or would it be more or less? do you have the 'present value' equation near by to calculate this?

OpenStudy (anonymous):

Makes sense. I'm just having trouble with the formula. Would it be A=2,000(1+0.08)^-19?

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

you have the right equation! But, we are trying to make the 'future' value of 6051.20 a 'present' value so replace the 2000 in your equation with 6051.20 A = F [(1 + i)^-n]

OpenStudy (anonymous):

So I'm trying to find the Principle (P)? as in 6051.20 = P(1+0.08)^-19 or include 2,000 as the (P)?

OpenStudy (anonymous):

ya!, the principle means 'present' or what the value would be today. so like this: P = 6051.20(1+0.08)^-19

OpenStudy (anonymous):

If so, the answer would be 26,115?

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

mmm... you have the right equation, but the math got mixed up P = 6051.20(1+0.08)^-19 P = 6051.20(1.08)^-19 [parenthesise first] P = 6051.20(0.2317) [then powers] P = 6051.20(0.2317) [then multipy] P = 1402.136

OpenStudy (anonymous):

THANK YOU SO MUCH!!

OpenStudy (anonymous):

^_^ your welcome